# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 4.7 Extension to the Fully Dynamic Setting
It is relatively straightforward to extend the algorithm from Section 4 to a setting where the input graph $G=(V,E)$ undergoes both edge insertions and edge deletions. As in Section 4, our fully dynamic algorithm works in phases, where each phase lasts for $\delta n$ updates in $G$.
@@
We maintain a maximal matching $M_{\texttt{core}}\subseteq E_{\texttt{core}}$ in $G_{\texttt{core}}$, using the decremental algorithm from Section 4. In a bit more detail, we handle an update in $G$ as follows. We first let the decremental algorithm from Section 4 handle the resulting update in $G_{\texttt{core}}$ and accordingly modify the matching $M_{\texttt{core}}$. Next, we scan through all the edges in $G_{\texttt{new}}$ and greedily add as many of these edges as we can on top of $M_{\texttt{core}}$, subject to the condition that the resulting edge-set $M_{\texttt{final}}\supseteq M_{\texttt{core}}$ remains a valid matching in $G$. This scan takes $O(|E_{\texttt{new}}|)=O(\delta n)$ time. It is easy to verify that the matching $M_{\texttt{final}}$ we obtain at the end of our scan is a maximal matching in $G$.
To summarize, with an additive overhead of $O(\delta n)$ update time, we can convert the decremental maximal matching algorithm from Section 4 into a fully dynamic algorithm. Finally, looking back at the analysis from Section 4.4, it is easy to verify that this additive overhead of $O(\delta n)$ does not degrade our overall asymptotic update time (the values of $\delta,B$ remain the same as in Section 4.4).
@@ 5.1 Structural Properties of Degree-Gap Graphs
Before proving Lemma 5.3, we show why it immediately implies Lemma 5.2.
Proof of Lemma 5.2.
@@
Therefore, Lemma 5.3(1) implies that every $M$-matched vertex $u_{\ell}\in L$ has an alternating path $P$ of length $O(\log(n)/\gamma)$ from $u$ that starts with an edge $e\notin M$ and ends at an $M$-free vertex.
For each $M$-free vertex $u_{\ell}\in L$, if $u_{\ell}$ has an $M$-free neighbor $v_{r}$, then $(u_{\ell},v_{r})$ is the desired path of length one. Otherwise, $u_{\ell}$ has a neighbor $u_{r}^{(2)}$ where $(u_{r}^{(2)},u_{\ell}^{(2)})\in M$. By the above argument, there exists an alternating path $P_{2}$ from $u_{\ell}^{(2)}$ to some $M$-free vertex $v_{r}$ of length $O(\log(n)/\gamma)$. So the concatenated path $P=(u_{\ell},u_{r}^{(2)})\circ P_{2}$ is the desired path of length $O(\log(n)/\gamma)$. This completes the proof of the first part of Lemma 5.2.
Lemma 5.3(2) analogously implies the “moreover” part of Lemma 5.2. The concrete algorithm description is as follows:
@@
This “random alternating walk” algorithm will arrive at an $M$-free vertex in $\Theta(\log(n)/\gamma^{2})$ steps by Lemma 5.3 and by the correspondence between paths in $G_{M}$ and alternating paths in $G$ as discussed above. By repeating the random alternating walks for $k=O(\log(n)/\gamma)$ times, the failure probability is at most $(1-\Omega(\gamma))^{k}<1/{\rm poly}(n)$. Since each vertex has access to a binary search tree of its incident edges, each random walk step can be implemented in $O(\log n)$ time. Hence, the total running time is $O(\log n)\times k\times\Theta(\log(n)/\gamma^{2})=\Theta(\log^{3}(n)/\gamma^{3})$. ∎
The rest of this section is for proving Lemma 5.3.
Preliminaries for Lemma 5.3.
@@
Proof of Lemma 5.3.
Ultimately, Lemma 5.3 follows from only that $G$ has a $\gamma$-degree-gap at $X$.
@@
Given Proposition 5.5, we are ready to prove Lemma 5.3 below.
Proof of Lemma 5.3(1).
@@
It is well-known that when $\Phi(G_{\mathrm{Eu}})\geq\gamma$, then the diameter of $G_{\mathrm{Eu}}$ is at most $O(\frac{\log(E(G_{\mathrm{Eu}}))}{\gamma})=O(\frac{\log n}{\gamma})$. So, by Proposition 5.5, for every non-sink vertex $u$, there exists a path $P_{\mathrm{Eu}}$ from $u$ to $t$ in $G_{\mathrm{Eu}}$ of length $O(\frac{\log n}{\gamma})$.
@@
Before proving Lemma 5.3(2), we review facts related to random walks on Eulerian expanders.
@@
In particular, after $k\geq O(\log(n)/\phi^{2})$ steps, any random walk end at $t$ with probability $\approx\frac{\mathrm{deg}_{G}(t)}{\mathrm{vol}_{G}(V)}$. The assumption about the self loops is a mild technical condition to ensure that the distribution of the random walks in $G$ is not periodic and does converge. Theorem 5.6 follows from the well-known facts in literature. We give its proof for completeness in Appendix A.
Proof of Lemma 5.3(2).
@@
First, we claim $\mathrm{deg}_{G_{\mathrm{Eu}}}(t)\geq\frac{\gamma}{2}\mathrm{vol}_{G_{\mathrm{Eu}}}(V(G_{\mathrm{Eu}}))$. To see this, consider $S=V(G_{\mathrm{Eu}})\setminus\{t\}$. We have $|E_{G_{M}}(S,t)|\geq\gamma\mathrm{vol}_{G_{M}}^{{\rm out}}(S)$ by Equation 5. The inequality follows because $\mathrm{deg}_{G_{\mathrm{Eu}}}(t)=\mathrm{deg}_{G_{M}}^{{\rm in}}(t)=|E_{G_{M}}(S,t)|$ and $\mathrm{vol}_{G_{\mathrm{Eu}}}(V(G_{\mathrm{Eu}}))=\mathrm{vol}_{G_{M}}^{{\rm out}}(S)+\mathrm{deg}_{G_{M}}^{{\rm out}}(t)\leq\mathrm{vol}_{G_{M}}^{{\rm out}}(S)+\mathrm{vol}_{G_{M}}^{{\rm in}}(S)\leq 2\mathrm{vol}_{G_{M}}^{{\rm out}}(S)$.
In order to apply Theorem 5.6, for each vertex $v$ in $G_{\mathrm{Eu}}$, add $\mathrm{deg}_{G_{\mathrm{Eu}}}(v)$ many self-loops at $v$. Let $G^{\prime}_{\mathrm{Eu}}$ be the resulting graph. We have $\Phi(G^{\prime}_{\mathrm{Eu}})\geq\Phi(G_{\mathrm{Eu}})/2\geq\gamma/2$ by Proposition 5.5. Let $k=100\log(n)/\gamma^{2}$. By applying Theorem 5.6 to $G^{\prime}_{\mathrm{Eu}}$, we have that, for every vertex $v$, a $k$-step random walk from $v$ in $G^{\prime}_{\mathrm{Eu}}$ will end at $t$ with probability at least
@@ 5.2 Approach 1: Random Augmenting Paths
In this section, we prove Lemma 5.1(2). The proof is quite simple given Lemma 5.2. First of all, we implement $\mathrm{Init}(M_{0})$ straightforwardly in $O(|M_{0}|)$ time.
@@
Thus, $(G^{\prime},M,u)$ is a valid input for Lemma 5.2 and, to implement $\mathrm{Augment}(u)$, we simply feed $(G^{\prime},M,u)$ to Lemma 5.2. It will then return an alternating path from $u$ in $G^{\prime}$ in $O(\log^{3}(n)/\gamma^{3})$ time with high probability. Then, we augment $M$ along this path, which matches $u$ as desired.
Although the algorithm is randomized, we use fresh randomness on each call to Lemma 5.2. So the correctness does not depend on the previous random choices. Thus, the algorithm works against an adaptive adversary.
@@ 5.3 Approach 2: Decremental Shortest Paths in Residual Graphs
Here, we prove Lemma 5.1(1). Using a standard reduction, the algorithm will maintain a graph $G_{\textrm{res}}$ with dummy sink $t$ such that for any free vertex $u$, a $ut$ path in $G_{\textrm{res}}$ corresponds to a $ut$ augmenting path in $G$. Thus, whenever the data structure encounters an $\textrm{Augment}(u)$ operation, it suffices to find a path from $u$ to $t$ in $G_{\textrm{res}}$. To make this last step efficient, the algorithm will make use of an existing data structure for dynamic shortest paths, which was first developed by Even and Shiloach [SE81] and later generalized to directed graphs by King [Kin99].
@@
Let $G$ be a dynamic directed graph with positive integer weights and a dedicated sink vertex $t$. Let $\mathrm{dist}(v,t)$ be the shorest $vt$ distance in the current version of $G$. There exists a data structure ES that maintains a shortest path tree $T$ to sink $t$ in $G$ [^6] Formally, $T$ is a tree on the vertex $V$ such that for any $v\in V$, the unique simple $vt$ path in $T$ is a shortest $vt$ path in $G$ where $G$ is subject to the following operations:
@@ 5.3.3 Pseudocode for LPM
[…]
@@ 6       Delete any previous ES data structure
7       Construct the graph $G_{\textrm{res}}$ from $G$ and $M$, setting all edge weights to $1$
@@ 35       Insert all edges of the reverse path $P^{\textrm{rev}}_{uv}$ into $G_{\textrm{res}}$
37       Delete all edges in $P{uv}$ from $G_{\textrm{res}}$
@@ 5.3.4 Analysis of LPM
[…]
@@
Consider any epoch. Whenever $\textrm{LPM}.\textrm{Delete}(u,v)$ is called, with $v\in R$, we say that $v$ is affected. Let $R_{\textrm{aff}}$ contain all affected right vertices and let $R_{\textrm{un}}=R\setminus R_{\textrm{aff}}$. We say that a left vertex $v\in L$ is affected if it contains fewer than $X(1-\gamma/2)$ neighbors in $L_{\textrm{un}}$ and unaffected otherwise: define $L_{\textrm{aff}}$ and $L_{\textrm{un}}$ accordingly. Define $G_{\textrm{un}}=G[L_{\textrm{un}},R_{\textrm{un}}]$. Finally, define $R^{M}_{\textrm{aff}}$ to contain all vertices in $R_{\textrm{aff}}$ and all right vertices that are matched to a vertex in $L_{\textrm{aff}}$; define $R^{M}_{\textrm{un}}=R\setminus R^{M}_{\textrm{aff}}$. When a new epoch begins (because ResetES is called), all vertices are again considered unaffected.
@@
All the following properties hold throughout an epoch:
  1. 1.
    We always have $|R_{\textrm{aff}}|\leq q_{\textrm{ep}}$
  2. 2.
    We always have $|L_{\textrm{aff}}|=O(|R_{\textrm{aff}}|/\gamma)$ and $R^{M}_{\textrm{aff}}=O(|R_{\textrm{aff}}|/\gamma)$.
  3. 3.
    The graph $G_{\textrm{un}}$ has $\gamma/2$-degree-gap at $X(1-\gamma/2)$
@@
(1) and (3) follow immediately from the definitions of $R_{\textrm{aff}}$ and $G_{\textrm{un}}$. For (2), recall by Invariant 5.12 that every $u\in L$ has $\mathrm{deg}(u)\geq X$, so $u$ can only be in $L_{\textrm{aff}}$ if it has at least $X\gamma/2$ neighbors in $R_{\textrm{aff}}$. The bound on $|L_{\textrm{aff}}|$ then follows from the fact that, by definition of degree gap, every vertex in $R_{\textrm{aff}}$ has degree at most $X(1-\gamma)<X$. The bound on $|R^{M}_{\textrm{aff}}|$ follows from the fact that $|R^{M}_{\textrm{aff}}|\leq|R_{\textrm{aff}}|+|L_{\textrm{aff}}|$. ∎
Claim 5.17.
Every vertex $v\in L_{\textrm{un}}\cup R^{M}_{\textrm{un}}$ has $\mathrm{dist}_{G_{\textrm{res}}}(v,t)=O(\log(n)/\gamma)$
Proof.
We prove the claim for $v\in L_{\textrm{un}}$. The proof for $v\in R^{M}_{\textrm{un}}$ immediately follows, because either $v$ is free and we can use the edge $(v,t)$ of weight $1$, or $v$ is matched to some $u\in L_{\textrm{un}}$, in which case we take the edge $(v,u)$ and then apply the claim from $u\in L$.
We thus assume $v\in L_{\textrm{un}}$. Since $G_{\textrm{un}}$ has $\gamma/2$-degree-gap at $X(1-\gamma/2)$, by Lemma 5.2 there exists an alternating path from $v$ to a free vertex $x\in R_{\textrm{un}}$ of length $O(\log(n)/\gamma)$. This corresponds to an $xv$ path of the same length in $G_{\textrm{res}}$ (Observation 5.9), which combined with the edge $(x,t)$ of weight $1$ (it has weight $1$ because $x\in R_{\textrm{un}}$) yields the desired $(v,t)$ path. ∎
@@
[…]
The base case is that while no vertices are affected, all weights in $G_{\textrm{res}}$ are $1$, so the subclaim follows from Lemma 5.2. Now assume that the subclaim holds for $|R_{\textrm{aff}}|\leq k-1$ and let $v_{k}\in R$ be the kth vertex in $R_{\textrm{aff}}$, which can only occur as a result of some $\textrm{LPM}.\textrm{Delete}(u,v_{k})$. The LPM data structure adds edge $(v_{k},t)$ to $G_{\textrm{res}}$; the weight of this edge is the distance $\mathrm{dist}(v_{k},t)$ before the deletion of $(u,v_{k})$. We now complete the proof by showing that before the deletion of $(u,v_{k})$, $\mathrm{dist}(v_{k},t)\leq kC\log(n)/\gamma$: by Lemma 5.2 there exists an augmenting path of length $\leq C\log(n)/\gamma$ in $G$ from $v_{k}$ to some free vertex $v^{\prime}\in R$, and by the induction hypothesis there is an edge $(v^{\prime},t)\in G_{\textrm{res}}$ of weight $w(v^{\prime},t)\leq(k-1)C\log(n)/\gamma$. ∎
@@
$\Delta\sum_{v\in V}\mathrm{dist}_{\textrm{max}}(v,t)$ (Δ is max $\leq\quad(\Delta\textrm{ is max degree})$
$\Delta\sum_{v\in L_{\textrm{un}}\cup R^{M}_{\textrm{un}}}\mathrm{dist}_{\textrm{max}}(v,t)+\Delta\sum_{v\in L_{\textrm{aff}}\cup R^{M}_{\textrm{aff}}}\mathrm{dist}_{\textrm{max}}(v,t)$$\leq$
@@
[…]
@@
[…]

### v2

@@ 4.7 Extension to the Fully Dynamic Setting
It is relatively straightforward to extend the algorithm from Section˜4 to a setting where the input graph $G=(V,E)$ undergoes both edge insertions and edge deletions. As in Section˜4, our fully dynamic algorithm works in phases, where each phase lasts for $\delta n$ updates in $G$.
@@
We maintain a maximal matching $M_{\texttt{core}}\subseteq E_{\texttt{core}}$ in $G_{\texttt{core}}$, using the decremental algorithm from Section˜4. In a bit more detail, we handle an update in $G$ as follows. We first let the decremental algorithm from Section˜4 handle the resulting update in $G_{\texttt{core}}$ and accordingly modify the matching $M_{\texttt{core}}$. Next, we scan through all the edges in $G_{\texttt{new}}$ and greedily add as many of these edges as we can on top of $M_{\texttt{core}}$, subject to the condition that the resulting edge-set $M_{\texttt{final}}\supseteq M_{\texttt{core}}$ remains a valid matching in $G$. This scan takes $O(|E_{\texttt{new}}|)=O(\delta n)$ time. It is easy to verify that the matching $M_{\texttt{final}}$ we obtain at the end of our scan is a maximal matching in $G$.
To summarize, with an additive overhead of $O(\delta n)$ update time, we can convert the decremental maximal matching algorithm from Section˜4 into a fully dynamic algorithm. Finally, looking back at the analysis from Section˜4.4, it is easy to verify that this additive overhead of $O(\delta n)$ does not degrade our overall asymptotic update time (the values of $\delta,B$ remain the same as in Section˜4.4).
@@ 5.1 Structural Properties of Degree-Gap Graphs
Before proving Lemma˜5.3, we show why it immediately implies Lemma˜5.2.
Proof of Lemma˜5.2.
@@
Therefore, Lemma˜5.3(1) implies that every $M$-matched vertex $u_{\ell}\in L$ has an alternating path $P$ of length $O(\log(n)/\gamma)$ from $u$ that starts with an edge $e\notin M$ and ends at an $M$-free vertex.
For each $M$-free vertex $u_{\ell}\in L$, if $u_{\ell}$ has an $M$-free neighbor $v_{r}$, then $(u_{\ell},v_{r})$ is the desired path of length one. Otherwise, $u_{\ell}$ has a neighbor $u_{r}^{(2)}$ where $(u_{r}^{(2)},u_{\ell}^{(2)})\in M$. By the above argument, there exists an alternating path $P_{2}$ from $u_{\ell}^{(2)}$ to some $M$-free vertex $v_{r}$ of length $O(\log(n)/\gamma)$. So the concatenated path $P=(u_{\ell},u_{r}^{(2)})\circ P_{2}$ is the desired path of length $O(\log(n)/\gamma)$. This completes the proof of the first part of Lemma˜5.2.
Lemma˜5.3(2) analogously implies the “moreover” part of Lemma˜5.2. The concrete algorithm description is as follows:
@@
This “random alternating walk” algorithm will arrive at an $M$-free vertex in $\Theta(\log(n)/\gamma^{2})$ steps by Lemma˜5.3 and by the correspondence between paths in $G_{M}$ and alternating paths in $G$ as discussed above. By repeating the random alternating walks for $k=O(\log(n)/\gamma)$ times, the failure probability is at most $(1-\Omega(\gamma))^{k}<1/{\rm poly}(n)$. Since each vertex has access to a binary search tree of its incident edges, each random walk step can be implemented in $O(\log n)$ time. Hence, the total running time is $O(\log n)\times k\times\Theta(\log(n)/\gamma^{2})=\Theta(\log^{3}(n)/\gamma^{3})$. ∎
The rest of this section is for proving Lemma˜5.3.
Preliminaries for Lemma˜5.3.
@@
Proof of Lemma˜5.3.
Ultimately, Lemma˜5.3 follows from only that $G$ has a $\gamma$-degree-gap at $X$.
@@
Given Proposition˜5.5, we are ready to prove Lemma˜5.3 below.
Proof of Lemma˜5.3(1).
@@
It is well-known that when $\Phi(G_{\mathrm{Eu}})\geq\gamma$, then the diameter of $G_{\mathrm{Eu}}$ is at most $O(\frac{\log(E(G_{\mathrm{Eu}}))}{\gamma})=O(\frac{\log n}{\gamma})$. So, by Proposition˜5.5, for every non-sink vertex $u$, there exists a path $P_{\mathrm{Eu}}$ from $u$ to $t$ in $G_{\mathrm{Eu}}$ of length $O(\frac{\log n}{\gamma})$.
@@
Before proving Lemma˜5.3(2), we review facts related to random walks on Eulerian expanders.
@@
In particular, after $k\geq O(\log(n)/\phi^{2})$ steps, any random walk end at $t$ with probability $\approx\frac{\mathrm{deg}_{G}(t)}{\mathrm{vol}_{G}(V)}$. The assumption about the self loops is a mild technical condition to ensure that the distribution of the random walks in $G$ is not periodic and does converge. Theorem˜5.6 follows from the well-known facts in literature. We give its proof for completeness in Appendix˜A.
Proof of Lemma˜5.3(2).
@@
First, we claim $\mathrm{deg}_{G_{\mathrm{Eu}}}(t)\geq\frac{\gamma}{2}\mathrm{vol}_{G_{\mathrm{Eu}}}(V(G_{\mathrm{Eu}}))$. To see this, consider $S=V(G_{\mathrm{Eu}})\setminus\{t\}$. We have $|E_{G_{M}}(S,t)|\geq\gamma\mathrm{vol}_{G_{M}}^{{\rm out}}(S)$ by Equation˜5. The inequality follows because $\mathrm{deg}_{G_{\mathrm{Eu}}}(t)=\mathrm{deg}_{G_{M}}^{{\rm in}}(t)=|E_{G_{M}}(S,t)|$ and $\mathrm{vol}_{G_{\mathrm{Eu}}}(V(G_{\mathrm{Eu}}))=\mathrm{vol}_{G_{M}}^{{\rm out}}(S)+\mathrm{deg}_{G_{M}}^{{\rm out}}(t)\leq\mathrm{vol}_{G_{M}}^{{\rm out}}(S)+\mathrm{vol}_{G_{M}}^{{\rm in}}(S)\leq 2\mathrm{vol}_{G_{M}}^{{\rm out}}(S)$.
In order to apply Theorem˜5.6, for each vertex $v$ in $G_{\mathrm{Eu}}$, add $\mathrm{deg}_{G_{\mathrm{Eu}}}(v)$ many self-loops at $v$. Let $G^{\prime}_{\mathrm{Eu}}$ be the resulting graph. We have $\Phi(G^{\prime}_{\mathrm{Eu}})\geq\Phi(G_{\mathrm{Eu}})/2\geq\gamma/2$ by Proposition˜5.5. Let $k=100\log(n)/\gamma^{2}$. By applying Theorem˜5.6 to $G^{\prime}_{\mathrm{Eu}}$, we have that, for every vertex $v$, a $k$-step random walk from $v$ in $G^{\prime}_{\mathrm{Eu}}$ will end at $t$ with probability at least
@@ 5.2 Approach 1: Random Augmenting Paths
In this section, we prove Lemma˜5.1(2). The proof is quite simple given Lemma˜5.2. First of all, we implement $\mathrm{Init}(M_{0})$ straightforwardly in $O(|M_{0}|)$ time.
@@
Thus, $(G^{\prime},M,u)$ is a valid input for Lemma˜5.2 and, to implement $\mathrm{Augment}(u)$, we simply feed $(G^{\prime},M,u)$ to Lemma˜5.2. It will then return an alternating path from $u$ in $G^{\prime}$ in $O(\log^{3}(n)/\gamma^{3})$ time with high probability. Then, we augment $M$ along this path, which matches $u$ as desired.
Although the algorithm is randomized, we use fresh randomness on each call to Lemma˜5.2. So the correctness does not depend on the previous random choices. Thus, the algorithm works against an adaptive adversary.
@@ 5.3 Approach 2: Decremental Shortest Paths in Residual Graphs
Here, we prove Lemma˜5.1(1). Using a standard reduction, the algorithm will maintain a graph $G_{\textrm{res}}$ with dummy sink $t$ such that for any free vertex $u$, a $ut$ path in $G_{\textrm{res}}$ corresponds to a $ut$ augmenting path in $G$. Thus, whenever the data structure encounters an $\textrm{Augment}(u)$ operation, it suffices to find a path from $u$ to $t$ in $G_{\textrm{res}}$. To make this last step efficient, the algorithm will make use of an existing data structure for dynamic shortest paths, which was first developed by Even and Shiloach [SE81] and later generalized to directed graphs by King [Kin99].
@@
Let $G$ be a dynamic directed graph with positive integer weights and a dedicated sink vertex $t$. Let $\mathrm{dist}(v,t)$ be the shorest $vt$ distance in the current version of $G$. There exists a data structure ES that maintains a shortest path tree $T$ to sink $t$ in GG [^6] Formally, $T$ is a tree on the vertex $V$ such that for any $v\in V$, the unique simple $vt$ path in $T$ is a shortest $vt$ path in $G$ where $G$ is subject to the following operations:
@@ 5.3.3 Pseudocode for LPM
We saw above that deletions will cause the weights of the ES tree to increase. As a result, the more deletions we have the more $\mathrm{dist}_{\textrm{max}}(v,t)$ will increase, and the less efficient the ES tree will become. In order to maintain efficiency, we introduce a new parameter $q_{\textrm{ep}}$ (ep for epoch), and every $q_{\textrm{ep}}$ calls to $\textrm{LPM}.\textrm{Delete}$ we reset all weights in $G_{\textrm{res}}$ to $1$ and build a new ES data structure from scratch.
@@ 6   Delete any previous ES data structure
7   Construct the graph $G_{\textrm{res}}$ from $G$ and $M$, setting all edge weights to 11
@@ 35   Insert all edges of the reverse path $P^{\textrm{rev}}_{uv}$ into $G_{\textrm{res}}$
37   Delete all edges in PuvP{uv} from $G_{\textrm{res}}$
@@ 5.3.4 Analysis of LPM
There are at most $q_{\textrm{ep}}$ calls to $\textrm{LPM}.\textrm{Delete}$ in a single epoch.
From now on, we will fix a single epoch and analyze the total runtime of LPM during that epoch.
@@
Consider any epoch. Whenever $\textrm{LPM}.\textrm{Delete}(u,v)$ is called, with $v\in R$, we say that $v$ is affected. Define $R_{\textrm{aff}}$ to contain all affected vertices.
@@
Throughout an epoch, we have $|R_{\textrm{aff}}|\leq q_{\textrm{ep}}$. Moreover, if $M$ is the current matching, then for any free vertex $v$, we have $w(v,t)>1$ only if $v\in R_{\textrm{aff}}$.
Structural Lemma
The key to the analysis is showing that within a single epoch, most vertices in $T_{\mathrm{res}}$ always have low distance to $t$. We cannot directly apply Lemma˜5.2 because edges in $T_{\mathrm{res}}$ can have positive weight. But by Observation˜5.16, the number of edges $(v,t)$ with weight >1>1 is at most $q_{\textrm{ep}}$. We are thus able to show that the majority of vertices are unaffected and have $\mathrm{dist}(v,t)=O(\log(n)/\gamma)$, as in Lemma˜5.2.
To formalize the above intuition, we use the following structural lemma, which is a generalization of Lemma˜5.2. It shows that in a degree-gap graph, if we mark an arbitrary small set of vertices in $R$, then most vertices in $L$ still have a short augmenting path to an unmarked free vertex.
Lemma 5.17.
Let $G=(L,R,E)$ be a bipartite graph with $\gamma$-degree-gap at $X$ and let $|L|=n$. Let $M$ be any left-perfect matching in $G$. Then, given any set of marked right-vertices $R_{\textrm{mark}}\subseteq R$, there exists a set $L_{\textrm{far}}\subseteq L$ with $|L_{\textrm{far}}|\leq\frac{2}{\gamma}\cdot|R_{\textrm{mark}}|$, such that for every vertex $u\in L\setminus L_{\textrm{far}}$, there exists an alternating path $P$ from $u$ that starts with an edge $e\notin M$, ends at an unmarked $M$-free vertex, and has length at most $4\ln(n)/\gamma=O(\log(n)/\gamma)$.
We defer the proof of Lemma˜5.17 to the next section. Here we show how to complete the analysis of LPM using the lemma as a black box.
Corollary 5.18.
Consider any epoch. There exists a set of vertices $V_{\textrm{far}}$ with $|V_{\textrm{far}}|=O(q_{\textrm{ep}}/\gamma)$, such that for every $v\in V\setminus V_{\textrm{far}}$, $\mathrm{dist}_{\textrm{max}}(v,t)=O(\log(n)/\gamma)$. [^7] Since the entire ES-tree data structure is reset at the end of the epoch, $\mathrm{dist}_{\textrm{max}}$ refers to the maximum distance within a single epoch.
@@
Proof of Corollary Let $G$ be the graph and $M$ the matching at end of the epoch. Recall that $G$ initially has $\gamma$-degree-gap at $X$ by the statement of Lemma 5.1. Observe that $G$ continues to have $\gamma$-degree-gap at $X$ throughout the execution of LPM: the degree constraint for $R$ holds automatically because the update sequence is decremental, and the degree constraint for $L$ holds because Line 1 of algorithm LPM removes violating vertices from the graph.
Applying Lemma˜5.17 with $R_{\textrm{mark}}=R_{\textrm{aff}}$ yields a set $L_{\textrm{far}}$ such that every $v\in L\setminus L_{\textrm{far}}$ has an alternating path of length $O(\log(n)/\gamma)$ to a free vertex $x\notin R_{\textrm{aff}}$. By observation Observation˜5.16, we have $w(x,t)=1$, so $\mathrm{dist}(v,t)=O(\log(n)/\gamma)$. Since distances in the ES-tree $T_{\mathrm{res}}$ are only increasing, and since we are considering the graph at the end of the epoch, we have $\mathrm{dist}_{\textrm{max}}(v,t)=\mathrm{dist}(v,t)=O(\log(n)/\gamma)$.
We now need to consider vertices in $R$ at the end of the epoch. Any $v\in R$ that is free and not in $R_{\textrm{aff}}$ has $\mathrm{dist}_{\textrm{max}}(v,t)=\mathrm{dist}(v,t)=w(v,t)=1$. Any vertex $v\in R$ that is matched to a vertex in $L\setminus L_{\textrm{far}}$ can follow a single edge to its mate and then an augmenting path to $t$, so also has $\mathrm{dist}_{\textrm{max}}(v,t)=\mathrm{dist}(v,t)=O(\log(n)/\gamma)$. This leaves at most $|R_{\textrm{aff}}|+|L_{\textrm{far}}|=O(q_{\textrm{ep}}/\gamma)$ vertices, all of which can be added to $V_{\textrm{far}}$. ∎
We couple the above corollary with a bound on $\mathrm{dist}_{\textrm{max}}(v,t)$ that applies even to $v\in V_{\textrm{far}}$.
@@
Recall that each $\textrm{LPM}.\textrm{Delete}$ adss at most one vertex to $R_{\textrm{aff}}$. We will prove the following subclaim for some large enough constant $C$: while $|R_{\textrm{aff}}|\leq k$, the maximum weight in $G_{\textrm{res}}$ is $\leq(k+1)C\frac{\log(n)}{\gamma}$; Claim˜5.19 then follows from ˜5.12 and Lemma˜5.2 (recall that only the edges into $t$ can have weight >1>1, so every $vt$ path contains at most one edge of weight >1>1).
The base case is that while no vertices are affected, all weights in $G_{\textrm{res}}$ are $1$, so the subclaim follows from Lemma˜5.2. Now assume that the subclaim holds for $|R_{\textrm{aff}}|\leq k-1$ and let $v_{k}\in R$ be the kth vertex in $R_{\textrm{aff}}$, which can only occur as a result of some $\textrm{LPM}.\textrm{Delete}(u,v_{k})$. The LPM data structure adds edge $(v_{k},t)$ to $G_{\textrm{res}}$; the weight of this edge is the distance $\mathrm{dist}(v_{k},t)$ before the deletion of $(u,v_{k})$. We now complete the proof by showing that before the deletion of $(u,v_{k})$, $\mathrm{dist}(v_{k},t)\leq kC\log(n)/\gamma$: by Lemma 5.2 there exists an augmenting path of length $\leq C\log(n)/\gamma$ in $G$ from $v_{k}$ to some free vertex $v^{\prime}\in R$, and by the induction hypothesis there is an edge $(v^{\prime},t)\in G_{\textrm{res}}$ of weight $w(v^{\prime},t)\leq(k-1)C\log(n)/\gamma$. ∎
@@
$O(\textrm{total \# updates}+\sum_{v\in V}\mathrm{deg}_{\textrm{max}}(v)\cdot\mathrm{dist}_{\textrm{max}}(v,t)).$
@@
$\Delta\sum_{v\in V}\mathrm{dist}_{\textrm{max}}(v,t)$ $\leq\quad(\Delta\textrm{ is max degree})$
$\Delta\sum_{v\in V\setminus V_{\textrm{far}}}\mathrm{dist}_{\textrm{max}}(v,t)+\Delta\sum_{v\in V_{\textrm{far}}}\mathrm{dist}_{\textrm{max}}(v,t)$ $\leq$
@@
We now need to bound the total number of updates to the ES tree. Every $\textrm{LPM}.\textrm{Delete}$ operation only causes $O(1)$ updates to the ES tree, for a total of $q_{\textrm{ep}}$. It remains to analyze $\textrm{LPM}.\textrm{Augment}(u)$ operations. Each such operation involves flipping every edge on the augmenting path, which amounts to $O(\mathrm{dist}(u,t))$ operations to the ES data structure, because $\mathrm{dist}(u,t)$ is an upper bound on the number of edges in the $ut$ augmenting path found by $\textrm{LPM}.\textrm{Augment}(u)$ (it may not be a tight upper bound because $G_{\textrm{res}}$ can have weights >1>1).
@@
$\textrm{[total \# updates] to }\textrm{{ES}}\leq\sum_{u\in L}\mathrm{dist}_{\textrm{max}}(u,t)+q_{\textrm{ep}}\cdot\frac{q_{\textrm{ep}}\log(n)}{\gamma}$